# Temporal Cross-Attention for Dynamic Embedding and Tokenization of Multimodal Electronic Health Records

### v1

@@ 2.1 Encoding dynamic temporality in clinical time series
Variable-specific encoding. A multivariate clinical time series often includes different categories of health variables (e.g., vital signs, laboratory tests, medications) that tend to exhibit distinct characteristics, numerical ranges, and temporal patterns. In recent EHR transformer implementations like BEHRT (Li et al., 2020)), all tokens within the sequence are embedded with the same encoder, which does not account for the heterogeneity among different variables. To address these existing challenges, our approach uses a separate encoder for each temporal clinical variable.
@@ 3.3 Models
Our embedding and tokenization scheme was designed for use with clinical transformers, and our primary classification model is a Longformer with sliding window attention. We compared our approach with several widely adopted baselines: (1) Tokenized GRUs using our embedding and tokenization scheme(Tan et al., 2020; Shi et al., 2021), (2) Tokenized XGBoost(Wang et al., 2020a; Liu et al., 2022), (3) BEHRT(Li et al., 2020), (4) Hi-BEHRT(Li et al., 2022a), (5) STraTS(Tipirneni & Reddy, 2022).
@@ 4.1 Time Series Modeling
Task Mean ICU AKI MV Mortality Wound Neurological Sepsis Cardiovascular VTE GRU + Attention 0.771±$\pm 0.04$ 0.857 0.718 0.783 0.816 0.712 0.753 0.791 0.762 0.747 XGBoost 0.765±$\pm 0.03$ 0.851 0.716 0.771 0.815 0.709 0.748 0.788 0.760 0.727 Transformer (BEHRT) 0.749±$\pm 0.01$ 0.843 0.701 0.765 0.800 0.701 0.725 0.770 0.748 0.699 Hi-BEHRT 0.781±$\pm 0.03$ 0.863 0.730 0.789 0.835 0.721 0.769 0.801 0.780 0.769 STraTS 0.797±$\pm 0.04$ 0.881 0.742 0.803 0.867 0.734 0.797 0.813 0.791 0.772 Transformer (Ours) 0.801±$\pm 0.02$ 0.883 0.749 0.810 0.853 0.739 0.800 0.811 0.797 0.774
Table 1 compares the area under the receiver operating characteristic curve (AUROC) for our benchmark multitask classification task. As shown in the table, our dynamic tokenization scheme-based transformer model outperform all baseline models with the highest mean AUROC of 0.801. We experimented with different types of variable-specific encoders (1-D convolutional layers (Kiranyaz et al., 2021) and transformer layers (Tipirneni & Reddy, 2022)) and found similar results. STraTS (Tipirneni & Reddy, 2022) slightly underperformed our approach, suggesting the utility of the added relative positional embeddings used by our approach. With the same tokenized sequence, GRU + Attention (mean AUROC: 0.771) outperformed transformer models with traditional tokenization scheme (mean AUROC: 0.749), indicating the advantages offered by our embedding and tokenization framework for other non-transformer modeling approaches.
@@ A.3 Data Preprocessing
Time Series Data For 14 intraoperative time series data, their variable names were converted to unique integer identifiers; the measured values for each variable was were normalized to zero mean and unit variance based on the values from the training set; their measurement time, in the format of “month/day/year hour:min:sec”, were first converted to unix timestamps and then also normalized similarly. For absolute positional indices, we assign one integer positional index for each token yet not enforcing the restriction that positional indices are unique and if different variables were measured at the same time. For relative positional embeddings, we generated the relative positional representation based on the GitHub code [^1] https://github.com/microsoft/MPNet/blob/master/pretraining/fairseq/modules/rel_multihead_attention.py for the original paper Shaw et al. (2018).
@@ A.4 Experiment Details
Longformer for processing long sequences Vanilla transformer models are largely limited by their capacity in processing long sequences (max_sequence = 512) because of its quadratic complexity with respect to the sequence length. Therefore, in this paper we used Longformer Beltagy et al. (2020) for time series modeling transformers. Longformer architecture introduces sliding window attention that allows the model to attend to only a subset of tokens within a window, reducing the overall computational complexity. This design makes Longformer well-suited for processing longer sequences (max_sequence = 4096).
@@ A.5 Related Work
Multimodal Representation Learning for Health Combining diverse sources of data sources in medical domain is promising for more comprehensive understanding of patients’ health conditions (Raghupathi & Raghupathi, 2014), more accurate health outcome predictions (Shickel & Bihorac, 2023), and building next-generation foundational medical models for generative AI (Moor et al., 2023). The core of this research effort is multimodal representation learning where all the modalities are projected to a common space while preserving information from the given modalities (Liang et al., 2022). Traditional data fusion methods, such as early and late fusion, are insufficient to learn the correlations and dependencies among different modalities (Lahat et al., 2015). Recently, transformer-based architecture, thanks to its superior ability to capture cross-modal interactions by self-attention and its variants (Xu et al., 2023), has achieved great success in various multimodal machine learning tasks in different domains, such as multimodal action recognition (Wang et al., 2020b), image segmentation (Xiao et al., 2023), and affect detection (Ma et al., 2023).

### v2

@@ 2.1 Encoding dynamic temporality in clinical time series
Variable-specific encoding. A multivariate clinical time series often includes different categories of health variables (e.g., vital signs, laboratory tests, medications) that tend to exhibit distinct characteristics, numerical ranges, and temporal patterns. In recent EHR transformer implementations like BEHRT (Li et al., 2020)), all tokens within the sequence are embedded with the same encoder, which does not account for the heterogeneity among different variables. To address these existing challenges, we propose to use a separate encoder for each clinical variable for intra-variable temporal dynamics, and then concatenate the outputs of the seperate encoders to learn the inter-variable correlation and dependencies.
@@ 3.3 Models
Our embedding and tokenization scheme was designed for use with clinical transformers, and our primary classification model is a Longformer with sliding window attention. We compared the model trained with our dynamic embedding and tokenization scheme with several widely adopted baselines:
Tokenized gated recurrent units (GRUs) with attention: the tokenized sequential data was provided as input to a multi-layer GRU, followed by a self-attention layer (Tan et al., 2020; Shi et al., 2021).
Tokenized XGBoost: an XGBoost model was trained on tokenized sequential data (Wang et al., 2020a; Liu et al., 2022).
BHERT(Li et al., 2020), a widely used baseline transformer model for EHR data. In this baseline model, we removed variable-specific encoders and used one single embedding layer to encode sequential data. We also removed the relative positional embedding and time embedding, as they were not included in BHERT tokenization scheme.
Hi-BEHRT (Li et al., 2022a), a hierarchical transformer-based model for EHR data, which employs a sliding window to partition the long sequence into smaller segments. Within each segment, a transformer was utilized as a local feature extractor to capture temporal dynamics.
Self-supervised Transformer for Time-Series (STraTS) (Tipirneni & Reddy, 2022). STraTS uses a unique transformer to encode each variable with transformers and then uses a self-attention layer to generate the time-series embedding.
@@ 4.1 Time Series Modeling
Task Mean ICU AKI MV Mortality Wound Neurological Sepsis Cardiovascular VTE GRU + Attention 0.771±$\pm 0.04$ 0.857 0.718 0.783 0.816 0.712 0.753 0.791 0.762 0.747 XGBoost 0.765±$\pm 0.03$ 0.851 0.716 0.771 0.815 0.709 0.748 0.788 0.760 0.727 Transformer (BEHRT) 0.749±$\pm 0.01$ 0.843 0.701 0.765 0.800 0.701 0.725 0.770 0.748 0.699 Hi-BEHRT 0.781±$\pm 0.03$ 0.863 0.730 0.789 0.835 0.721 0.769 0.801 0.780 0.769 STraTS 0.797±$\pm 0.04$ 0.881 0.742 0.803 0.857 0.734 0.797 0.813 0.791 0.772 Transformer (Ours) 0.801±$\pm 0.02$ 0.883 0.749 0.810 0.853 0.739 0.800 0.811 0.797 0.774
Table 1 compares the area under the receiver operating characteristic curve (AUROC) for our benchmark multitask classification task. As shown in the table, our dynamic tokenization scheme-based transformer model outperform all baseline models with the highest mean AUROC of 0.801. We experimented with different types of variable-specific encoders (1-D convolutional layers (Kiranyaz et al., 2021) and transformer layers (Tipirneni & Reddy, 2022)) and found similar results (For 1-D CNN, transformer, and linear, the mean AUROC score was 0.796, 0.800, and 0.798, respectively). STraTS (Tipirneni & Reddy, 2022) slightly underperformed our approach, suggesting the utility of the added relative positional embeddings used by our approach. With the same tokenized sequence, GRU + Attention (mean AUROC: 0.771) outperformed transformer models with traditional tokenization scheme (mean AUROC: 0.749), indicating the advantages offered by our embedding and tokenization framework for other non-transformer modeling approaches.
@@ A.3 Data Preprocessing
Time Series Data For 14 intraoperative time series data, their variable names were converted to unique integer identifiers; the measured values for each variable was were normalized to zero mean and unit variance based on the values from the training set; their measurement time, in the format of “month/day/year hour:min:sec”, were first converted to unix timestamps and then also normalized similarly. For absolute positional indices, we assign one integer positional index for each token yet not enforcing the restriction that positional indices are unique and if different variables were measured at the same time. For relative positional embeddings, we generated the relative positional representation based on the GitHub code [^1] https://github.com/microsoft/MPNet/blob/master/pretraining/fairseq/modules/rel_multihead_attention.py for the original paper (Shaw et al., 2018).
@@ A.4 Experiment Details
Longformer for processing long sequences Vanilla transformer models are largely limited by their capacity in processing long sequences (max_sequence = 512) because of its quadratic complexity with respect to the sequence length. Therefore, in this paper we used Longformer (Beltagy et al., 2020) for time series modeling transformers. Longformer architecture introduces sliding window attention that allows the model to attend to only a subset of tokens within a window, reducing the overall computational complexity. This design makes Longformer well-suited for processing longer sequences (max_sequence = 4096).
@@ A.5 Related Work
Multimodal Representation Learning for Health Combining diverse sources of data sources in medical domain is promising for more comprehensive understanding of patients’ health conditions (Raghupathi & Raghupathi, 2014), more accurate health outcome predictions (Shickel & Bihorac, 2023), and building next-generation foundational medical models for generative AI (Moor et al., 2023). The core of this research effort is multimodal representation learning where all the modalities are projected to a common space while preserving information from the given modalities (Liang et al., 2022). Traditional data fusion methods, such as early and late fusion, are insufficient to learn the correlations and dependencies among different modalities (Ma et al., 2022). Recently, transformer-based architecture, thanks to its superior ability to capture cross-modal interactions by self-attention and its variants (Xu et al., 2023), has achieved great success in various multimodal machine learning tasks in different domains, such as multimodal action recognition (Wang et al., 2020b), image segmentation (Xiao et al., 2023), and affect detection (Ma et al., 2023).